# Surrogate normal-forms for the numerical bifurcation and stability analysis of Navier-Stokes flows via machine learning

### v1

@@ 2.5 Numerical bifurcation analysis
$N(\mathbf{a},Re)=\mathbf{c}\cdot\left(\mathbf{a}-\mathbf{a}_{1}\right)+b(Re-Re_{1})=0,\qquad\mathbf{c}\equiv\dfrac{\left(\mathbf{a}_{1}-\mathbf{a}_{0}\right)^{\top}}{\delta s},\qquad b\equiv\dfrac{Re_{1}-Re_{0}}{\delta s}$ (17)
@@
[…]
@@ 4.4 Symmetry-breaking bifurcation
Once the performances of the reduced-order GPR model have been assessed by forecasting the flow dynamics for new conditions, the numerical bifurcation diagram is first computed by applying to the system (15) the continuation method described in Section 2.5. As a result, we obtain the bifurcation diagram shown in Fig. 14(b), which identifies the existence of three steady solutions — two stable (continuous red curves) and one unstable (dashed red curve) — for $Re>43.12$. For $Re<43.12$, a unique stable steady solution exists. However, the flip symmetry of the generic pitchfork bifurcation is clearly retrieved in the computed diagram. Due to the intrinsic systematic error of the learned model, we have indeed identified a perturbed pitchfork, in which the lower branch remains permanently stable, and the upper branch exhibits a turning point where a stable and an unstable solution meet. As recently shown by [100] in the context of residual neural network models of agent-based systems, a simple odd symmetry transformation of Eq. (15) is sufficient to recover the pitchfork bifurcation. In particular, given the identified right-hand side $g(Re,a_{1})$ of the Gaussian Process regression model, we compute
@@
Finally, by addressing the pre-image problem, the POD-GPR reduced-order solutions can be lifted back to the original high-dimensional physical space to reconstruct the full bifurcation diagram, which is shown in Fig. 14(d). For comparison, we have also reported the Navier-Stokes solutions (black dots in Fig. 14(d)) in terms of transverse velocity $v$ stored at the spatial location $x=20$, $y=0$. It can be seen that the high-dimensional bifurcation diagram is already well reconstructed by the minimal $P=1$ reduced-order model (black curve in Fig. 14(d)) in the whole Reynolds number range considered, with a maximum relative percentage error equal to 22% for $Re=70$. The accuracy of the reconstruction improves by reducing the maximum error down to 3% if one takes into account the leading $P=3$ reduced-order coordinates in the surrogate POD-GPR model (blue curve in Fig. 14(d)).
@@ 5 Conclusions
The potential of the proposed pipeline has been demonstrated on well-studied benchmark Navier-Stokes simulations over a parameter range, each featuring a single bifurcation type, for which POD-based reductions have proven adequate. A natural next step is to showcase the pipeline’s efficiency in constructing global normal-form ROMs capable of capturing the full dynamics, including turbulent regimes, across broader parameter domains. The concept for the construction of “global” surrogate normal-forms via machine learning, but for low-dimensional systems of ODEs, can be traced back in the 2000s [101] using explainable machine learning. To deal with the high-dimensionality of the spatio-temporal Navier-Stokes equations, this approach can be extended by incorporating the invariant manifold theory [11, 12]. Such an approach has been recently proposed in [49] for dissipative 1D PDEs (the Chafee-Infante reaction-diffusion PDE and the Kuramoto-Sivashinsky PDE) using autoencoders, but also POD and Diffusion Maps, to parametrize an approximate inertial manifold (AIM) resulting in a “global” surrogate normal-form of ODEs. Its potential for application to the Navier-Stokes equations is strongly indicated by our recently results in [42], where we have shown that while the POD fails to identify the actual intrinsic dimension of the underlying latent space (i.e., the correct dimension of the normal-form model), Diffusion Maps succeed in identifying the appropriate dimension of the global dynamics.
Acknowledgements. The authors acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. The PNRR MUR, projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM) partially supported C.S.

### v2

@@ 2.5 Numerical bifurcation analysis
$N(\mathbf{a},Re)=\mathbf{c}\cdot\left(\mathbf{a}-\mathbf{a}_{1}\right)+b(Re-Re_{1})=0,\qquad\mathbf{c}\equiv\dfrac{\left(\mathbf{a}_{1}-\mathbf{a}_{0}\right)^{\top}}{\delta s},\qquad b\equiv\dfrac{Re_{1}-Re_{0}}{\delta s},$ (18)
@@
$\left\{\begin{aligned} \dfrac{d\mathbf{a}}{d\tau}-T\mathbf{g}(\mathbf{a},Re)&=0,\\ \mathbf{a}(0)-\mathbf{a}(1)&=0,\end{aligned}\right$. (20)
@@ 4.3 Andronov-Hopf bifurcation
Figure 13: Numerical bifurcation diagram of the cylinder wake flow reduced-order model by increasing the Reynolds number ReRe (panel (a)). A supercritical Hopf bifurcation occurs at $Re_{cr}=48.48$: the stable fixed point solution ($a_{1}=a_{2}=0$) becomes unstable, giving rise to a branch of limit cycles LC (blue curve). In panel (b), the high-dimensional bifurcation diagram (black continuous curve) is reconstructed in terms of $v$ velocity component at the spatial location $x=10$, $y=0$. Values from the Navier-Stokes solutions are also reported for comparison (black dots). Panel (c): a complex-conjugate pair of eigenvalues (red circles) becomes unstable (i.e., Re($\sigma_{i})>0$ for $i=1,2$) for $Re>Re_{cr}$. Panel (d): limit cycle period $T$ (black curve) and Floquet multipliers $f_{1}$ and $f_{2}$ as a function of ReRe obtained from the limit cycles branch continuation.
@@ 4.4 Symmetry-breaking bifurcation
Once the performances of the reduced-order GPR model have been assessed by forecasting the flow dynamics for new conditions, the numerical bifurcation diagram is first computed by applying to the system (16) the continuation method described in Section 2.5. As a result, we obtain the bifurcation diagram shown in Fig. 14(b), which identifies the existence of three steady solutions — two stable (continuous red curves) and one unstable (dashed red curve) — for $Re>43.12$. For $Re<43.12$, a unique stable steady solution exists. However, the flip symmetry of the generic pitchfork bifurcation is clearly retrieved in the computed diagram. Due to the intrinsic systematic error, which is inherent to any learned surrogate model, the result is a perturbed pitchfork: the lower branch remains permanently stable, while the upper branch exhibits a turning point where a stable and an unstable solution meet. As recently shown by [77] in the context of residual neural network models of agent-based systems, applying a simple odd symmetry transformation to Eq. (16) is sufficient to correctly pin the pitchfork bifurcation. In particular, given the identified right-hand side $g(Re,a_{1})$ of the Gaussian Process regression model, we compute
@@
Finally, by addressing the pre-image problem, the POD-GPR reduced-order solutions can be lifted back to the original high-dimensional physical space to reconstruct the full bifurcation diagram, which is shown in Fig. 14(d). For comparison, we have also reported the Navier-Stokes solutions (black dots in Fig. 14(d)) in terms of transverse velocity $v$ stored at the spatial location $x=20$, $y=0$. It can be seen that the high-dimensional bifurcation diagram is already well reconstructed by the minimal $P=1$ reduced-order model (black curve in Fig. 14(d)) in the whole Reynolds number range considered, with a maximum relative percentage error equal to 22% for $Re=70$. This is consistent with the expectation that the accuracy of reduced-order surrogates decreases away from the bifurcation point, as normal-form representations are reliable mainly in its vicinity, with deviations growing at larger distances. Nevertheless, the accuracy of the reconstruction improves, reducing the maximum error to 3% when including the leading $P=3$ reduced-order coordinates in the surrogate POD-GPR model (blue curve in Fig. 14(d)).
@@ 5 Conclusions
The potential of the proposed pipeline has been demonstrated on well-studied benchmark Navier-Stokes simulations over a parameter range, each featuring a single bifurcation type, for which POD-based reductions have proven adequate. A natural next step is to showcase the pipeline’s efficiency in constructing global normal-form ROMs capable of capturing the full dynamics, including turbulent regimes, across broader parameter domains. The concept for the construction of “global” surrogate normal-forms via machine learning, but for low-dimensional systems of ODEs, can be traced back in the 2000s [100] using explainable machine learning. To deal with the high-dimensionality of the spatio-temporal Navier-Stokes equations, this approach can be extended by incorporating the invariant manifold theory [11, 12]. Such an approach has been recently proposed in [49] for dissipative 1D PDEs (the Chafee-Infante reaction-diffusion PDE and the Kuramoto-Sivashinsky PDE) using autoencoders, but also POD and Diffusion Maps, to parametrize an approximate inertial manifold (AIM) resulting in a “global” surrogate normal-form of ODEs. Its potential for application to the Navier-Stokes equations is strongly indicated by our recent results in [42], where we have shown that while the POD fails to identify the actual intrinsic dimension of the underlying latent space (i.e., the correct dimension of the normal-form model), Diffusion Maps succeed in identifying the appropriate dimension of the global dynamics.
Acknowledgements. A.D.P. acknowledges the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. C. S. acknowledges partial support from the PNRR MUR, projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).